# GaussianArt: Unified Modeling of Geometry and Motion for Articulated Objects

### v1

@@ 1 Introduction
Reconstructing articulated objects plays a central role in creating digital twins for robotic simulation and generic interaction modeling [36, 48, 13]. While recent progress [17, 33, 65, 54, 12, 38, 30, 79] has been made, most existing pipelines [65, 38, 30, 79] adopt a decoupled design as shown in Fig. 1: they first reconstruct geometry from two static observations and subsequently infer motion through part-wise alignment. This separation not only introduces redundant modeling and brittle optimization, but more critically, breaks the physical consistency across object states—the same region of geometry is reconstructed independently per state without a coherent articulation structure.
@@
As shown in Fig. 1 (top-left and top-middle), ArtGS [38] exemplifies this limitation: it separately models geometry and motion using per-state 3D Gaussians, then relies on clustering to estimate part motion. Such pipelines often suffer from incorrect part grouping and axis misalignment, especially when handling more than 2–3 moving parts.
@@
Several recent systems, including DigitalTwinArt [65] and ArtGS [38], attempt to address multi-part articulation by leveraging learned part-field or clustering heuristics. However, they are constrained by: (1) separate optimization of geometry and motion; (2) brittle initialization, often failing to produce reliable part decomposition under complex configurations; and (3) limited scalability, rarely tested beyond a dozen objects, and often constrained to two-part motion. Moreover, current benchmarks are narrow in scope—most evaluate on fewer than 20 objects with constrained topology and articulation patterns.
@@
We introduce GaussianArt, a physically consistent and scalable framework for articulated object reconstruction. Unlike prior methods, GaussianArt adopts a unified representation based on articulated 3D Gaussian primitives, where each Gaussian simultaneously encodes its part affiliation (via learned soft assignments) and its rigid motion (as a mixture of motion bases). This formulation enables geometry and motion to be co-optimized within a single differentiable structure, ensuring cross-state consistency and interpretability, as visualized in Fig. 1 (middle row).
@@
To evaluate scalability, we construct MPArt-90, a benchmark of 90 articulated objects across 20 categories, each with observations in two states and full articulation ground truth. As shown in Fig. 1 (top-right), GaussianArt correctly recovers fine-grained part structure and motion parameters, even with 7 components. In contrast, ArtGS suffers from cluster collapse, resulting in incorrect part grouping and misaligned joints. In terms of quality, Fig. 1 (bottom row) highlights the significant gains in both segmentation accuracy and reconstruction fidelity. GaussianArt reduces the dynamic part Chamfer Distance (CD) from 120.15 to 0.16, demonstrating an order-of-magnitude improvement.
@@ 2 Related Work
3D Gaussian Splatting (3DGS) [21, 51, 73] represents 3D scenes using ellipsoids with Gaussian distributions as geometric primitives. This representation enables novel view rendering through differentiable rasterization, making it highly efficient for both training and rendering. The superior performance of 3DGS has inspired a series of studies that apply it to more complex dynamic scenes, exploring various motion modeling approaches [74, 75, 15, 32, 59, 71]. Several studies, such as GART [25], utilize articulated motion primitives for modeling rigid human body movement via linear blend skinning (LBS), while Shape-of-Motion [62] applies a similar model for reconstructing dynamic scenes from monocular videos. Other approaches, like SC-GS [15], use sparsely distributed control points with motion priors, allowing user editing of motion, and Gaussian-Flow [32] employs a dual-domain deformation model considering frequency-domain transformations. However, most of these methods lack motion constraints and struggle to accurately estimate part-level articulated motion parameters.
@@
With the advancements of shape and dynamic modeling [29, 18, 40, 14, 56, 57, 85], many works begin leveraging 3D point clouds [77, 63, 28, 72, 64, 10, 11, 24, 84, 83], or images and videos (sometimes with depth) [16, 49, 19, 10, 53, 4, 42, 35, 27, 82, 69, 68, 41, 1, 67, 46, 9] to model articulated objects. However, the most effective way is to reconstruct high-fidelity digital twins through multi-state observation [47, 17, 55, 33, 65, 54, 6, 38, 66, 12, 30, 60, 79, 22].
Former works [47, 17] use different scene representations to reconstruct articulated objects. The emergence of neural radiance fields [44, 61, 37, 34, 81] and their high rendering fidelity have led to a series of works adopting neural implicit scene representations for articulated object reconstruction. [55, 33] are among the first efforts in reconstructing articulated objects based on neural rendering. Although promising results have been achieved for two-part objects, extending these approaches to multi-part scenarios remains challenging and lacks generalizability. Recent studies [65, 54, 6] have improved the accuracy of motion parameter prediction and demonstrated some capability in multi-part object reconstruction, but they struggle to generalize to objects with more parts and more complex motion combinations. ArtGS [38], built upon 3DGS, achieves certain improvements on multi-part objects, but its results are highly unstable and show poor generalization in large-scale evaluations. In this work, we propose a well-designed unified scalable reconstruction pipeline based on articulated Gaussians.
@@ 3.1 Preliminaries
3DGS [21] represents a 3D scene by a set of Gaussian Primitives, each defined as:
@@
To reconstruct articulated objects with multiple 1-DoF motion parts, we follow the two-state observation setting in DigitalTwinArt [65], requiring posed RGB-D sequences of the same scene in two states as input: , , where denotes the number of images, the RGB image, the depth map, the camera extrinsics, and the camera intrinsics..
@@ 3.3 Soft-to-hard Training Paradigm
Sometimes, part segmentation alone cannot fully constrain motion parameter learning, making extreme movements and occlusions difficult to model. To refine motion learning, we apply a feature-matching-based trajectory regularization term during hard training. For image from view at state , we select the top closest view from state and apply an image feature matching model [52] to establish 2D pixel correspondences , where denotes the number of matches and are pixel coordinates. We lift them to 3D space as using depth and camera parameters. Then, we apply a 3D locality filter on 3D maches to mitigate false correspondences, yielding a refined set of matching pairs . The transformation at the matching positions can be derived from Eq. 6. The trajectory regularization term is the average difference between the position obtained through transformation and that derived from the filtered correspondences (Fig. 3(b)):
@@ 3.4 Optimization
where , is the D-SSIM loss [21], and . We also use the segmentation loss:
@@ 3.5 Initialization
Part segmentation. To initialize and regularize Gaussian part weights , we leverage SAM2 [50], a foundation segmentation model pretrained on large-scale data. Since its zero-shot results often show inconsistent granularity on articulated objects, we fine-tune it on multi-view images and masks rendered from Partnet-Mobility (PM) [70], yielding a specialized model, Art-SAM. We then apply cross-view propagation to ensure multi-view consistency (see Supplementary Material Secs. 9 and 10). This method is robust and generalizable, as Art-SAM requires only light post-training to adapt to novel object categories or configurations.
@@ 4.1 Benchmark Data Generation
Articulated objects are mainly constructed from the PM dataset [70], from which we select 87 objects based on the diversity of categories, part number, and appearances. We use Blender with a procedural rendering pipeline [5, 7] to render multi-view images of the 3D models in the base assets. To increase the diversity of object states while providing adequate observation of the interior parts of the objects, we set the 1-DoF part-level motion parameter to two random states: the starting state lies between and the ending state is within (here denotes the ”fully-closed” state and denotes the ”fully-open” state). For the object at each motion state, we place the camera in a spherical region around the object and randomly sample 100 views for training and 20 views for testing, all at a resolution of .
Due to the limited availability of high-quality real-world articulated object data and the significant ground-truth errors caused by the inability to annotate the internal structures of real objects, we selected only three well-annotated real objects from the Multiscan [43] dataset. See Supplementary Material Sec. 18 for more results.
[…]
@@ 5.1 Implementations
We first evaluate several methods on several traditional datasets, and finally select our method and ArtGS [38] for scale-up evaluation on our MPArt-90 benchmark.
@@ 5.3 Ablation Studies
Part assignment. We evaluate our model under 3 different scenarios: (1) without part segmentation masks, (2) without part initialization, and (3) replacing the part assignment module with an MLP, as used in GART [25]. As shown in Table 2, without segmentation masks, Gaussians struggle to capture the complex motion of multiple parts. Additionally, relying solely on part segmentation supervision—without explicit part initialization—can negatively impact the geometric reconstruction of movable parts. When using MLPs for part assignment, the model fails entirely to learn part segmentation and motion for multi-part objects.
@@ 6.2 HSI
Furthermore, our Articulated Gaussians can be leveraged to generate 4D assets, enabling the modeling of HSI in dynamic environments. As depicted in Fig. 7, given Gaussian-based representations of humans, objects, and scenes as input, inspired by ZeroHSI [26], we could generate high-fidelity 4D scenes. We first render a static frame, which is then used as input to a text-guided video diffusion model to synthesize a corresponding video. This generated video serves as a supervisory signal to optimize the kinematic parameters of humans and objects, effectively lifting the original 3D assets into a coherent 4D representation.
Specifically, we employ SMPL [39] for human kinematic modeling and use our reconstructed digital twins, with articulation parameters, to represent objects with kinematics. Through a distillation process, the motion dynamics captured in the video are transferred to the 4D scene.
@@ 9 Art-SAM Training
To provide supervision for accurate part-level segmentation, we adopt an image segmentation model to generate segmentation masks for each view and then use multi-view reprojection consistency to propagate monocular masks across views. To this end, an image segmentation model is required. Visual foundation models [23, 50] have emerged as powerful tools for class-agnostic zero-shot or prompt-based image segmentation. However, when applied to articulated objects, the problems of over-segmentation and part ambiguity are common. Applying the models in a zero-shot manner often results in masks of erroneous granularity, where some masks cover overly fine-grained areas, while others fail to cover the full range of a complete interactable part (see the ’Zero-shot’ results of Fig. 10). Therefore, we fine-tuned SAM-v2.1 model [50] to generate segmentation masks for articulated objects at the desired granularity. The fine-tuned model can generate masks of the desired part-level granularity on synthetic objects (Fig. 10) and real objects (Fig. 11). This section covers the details of data preparation and the fine-tuning process.
Base Assets. We use the PM dataset [70] to create a dataset for fine-tuning SAM-v2.1. The dataset contains URDF-format models of articulated objects with motion parameters. We select 207 objects from 20 categories to generate pairs of images and corresponding segmentation masks for fine-tuning. For real-world assets, we select 50 objects from Multiscan [43].
Image Rendering. We use Blender with a procedural rendering pipeline [5, 7] to render multi-view images of the 3D models in the base assets. To increase the diversity of object states while providing adequate observation of the interior parts of the objects, we set the 1-DoF part-level motion parameter to random states between (here denotes the ”fully closed” state and denotes the ”fully open” state). For each object, we place the camera on a spherical region around the object and randomly sample 100 to 300 views, generating 130,000 training images at three resolution levels: , , and . Examples of training data are shown in Fig. 8.
Segmentation Map Rendering. To provide supervision at the desired granularity, we also adopt the procedural rendering pipeline with Blender [7], which allows the rendering of object masks. A training sample consists of an RGB image paired with its corresponding segmentation masks.
Training Setup. Starting from the pretrained SAM2.1-Hiera-B+ model, we fine-tune all model parameters for 40 epochs on the dataset we created from PM. The loss function is a weighted linear combination of MSE IoU loss, Dice loss [45], and focal loss [31]. In each epoch, the learning rate starts at and decays to using a cosine annealing scheduler.

### v2

@@ 1 Introduction
Reconstructing articulated objects plays a central role in creating digital twins for robotic simulation and generic interaction modeling [liu2024survey, pejic2022articulated, hu2018functionality]. While recent progress [Jiang_2022_ditto, liu2023paris, weng2024digitaltwin, swaminathan24leia, guo2025articulatedgs, liu2025artgs, lin2025splart, yu2025part] has been made, most existing pipelines [weng2024digitaltwin, liu2025artgs, lin2025splart, yu2025part] adopt a decoupled design as shown in LABEL:fig:teaser: they first reconstruct geometry from two static observations and subsequently infer motion through part-wise alignment. This separation not only introduces redundant modeling and brittle optimization, but more critically, breaks the physical consistency across object states—the same region of geometry is reconstructed independently per state without a coherent articulation structure.
@@
As shown in LABEL:fig:teaser (top-left and top-middle), ArtGS [liu2025artgs] exemplifies this limitation: it separately models geometry and motion using per-state 3D Gaussians, then relies on clustering to estimate part motion. Such pipelines often suffer from incorrect part grouping and axis misalignment, especially when handling more than 2–3 moving parts.
@@
Several recent systems, including DigitalTwinArt [weng2024digitaltwin] and ArtGS [liu2025artgs], attempt to address multi-part articulation by leveraging learned part-field or clustering heuristics. However, they are constrained by: (1) separate optimization of geometry and motion; (2) brittle initialization, often failing to produce reliable part decomposition under complex configurations; and (3) limited scalability, rarely tested beyond a dozen objects, and often constrained to two-part motion. Moreover, current benchmarks are narrow in scope—most evaluate on fewer than 20 objects with constrained topology and articulation patterns.
@@
We introduce GaussianArt, a physically consistent and scalable framework for articulated object reconstruction. Unlike prior methods, GaussianArt adopts a unified representation based on articulated 3D Gaussian primitives, where each Gaussian simultaneously encodes its part affiliation (via learned soft assignments) and its rigid motion (as a mixture of motion bases). This formulation enables geometry and motion to be co-optimized within a single differentiable structure, ensuring cross-state consistency and interpretability, as visualized in LABEL:fig:teaser (middle row).
@@
To evaluate scalability, we construct MPArt-90, a benchmark of 90 articulated objects across 20 categories, each with observations in two states and full articulation ground truth. As shown in LABEL:fig:teaser (top-right), GaussianArt correctly recovers fine-grained part structure and motion parameters, even with 7 components. In contrast, ArtGS suffers from cluster collapse, resulting in incorrect part grouping and misaligned joints. In terms of quality, LABEL:fig:teaser (bottom row) highlights the significant gains in both segmentation accuracy and reconstruction fidelity. GaussianArt reduces the dynamic part Chamfer Distance (CD) from 120.15 to 0.16, demonstrating an order-of-magnitude improvement.
@@ 2 Related Work
3D Gaussian Splatting (3DGS) [kerbl20233dgs, song2024sa, yang2024spectrally] represents 3D scenes using ellipsoids with Gaussian distributions as geometric primitives. This representation enables novel view rendering through differentiable rasterization, making it highly efficient for both training and rendering. The superior performance of 3DGS has inspired a series of studies that apply it to more complex dynamic scenes, exploring various motion modeling approaches [Yang_2024_Deformable, yang2023gs4d, Huang_2024_SCGS, Lin_2024_Gaussian_Flow, wan2024templatefree, xu2024longvolcap]. Several studies, such as GART [lei2024gart], utilize articulated motion primitives for modeling rigid human body movement via linear blend skinning (LBS), while Shape-of-Motion [som2024] applies a similar model for reconstructing dynamic scenes from monocular videos. Other approaches, like SC-GS [Huang_2024_SCGS], use sparsely distributed control points with motion priors, allowing user editing of motion, and Gaussian-Flow [Lin_2024_Gaussian_Flow] employs a dual-domain deformation model considering frequency-domain transformations. However, most of these methods lack motion constraints and struggle to accurately estimate part-level articulated motion parameters.
@@
With the advancements of shape and dynamic modeling [li2025megasam, jones2020shapeassembly, lu2023jigsaw, huang2011joint, tulsiani2016learning, tulsiani2017learning, zou20173d], many works begin leveraging 3D point clouds [Yi_2018_Deep_Part, wang2019shape2motion, Li_2020_Category_Level, yan2020rpm, Weng_2021_Captra, geng2023sage, geng2023gapartnet, kreber2025guiding, zhong20233d, zhong2022snake], or images and videos (sometimes with depth) [Jiang_2022_Opd, Qian_2022_understanding_3d_internet, kawana_2023_detection_based_part_level, geng2023sage, Sun_2024_OPDMulti, chen2024urdformer, mandi2024real2code, liu2024singapo, Li_2025_LocateNRotate, zhang2025iaao, xia2025drawer, wu2025dipo, lu2025dreamart, cao2025physx, wu2025predict, mo2021where2act, gao2025partrm] to model articulated objects. However, the most effective way is to reconstruct high-fidelity digital twins through multi-state observation [Mu_2021_ASDF, Jiang_2022_ditto, Tseng_2021_CLA-NeRF, liu2023paris, weng2024digitaltwin, swaminathan24leia, deng2024articulate, liu2025artgs, wu2025reartgs, guo2025articulatedgs, lin2025splart, wang2025self, yu2025part, kim2025screwsplat].
Former works [Mu_2021_ASDF, Jiang_2022_ditto] use different scene representations to reconstruct articulated objects. The emergence of neural radiance fields [mildenhall_2020_nerf, wang2021neus, liu2020neural, liu2024rip, yuan2024slimmerf] and their high rendering fidelity have led to a series of works adopting neural implicit scene representations for articulated object reconstruction. [Tseng_2021_CLA-NeRF, liu2023paris] are among the first efforts in reconstructing articulated objects based on neural rendering. Although promising results have been achieved for two-part objects, extending these approaches to multi-part scenarios remains challenging and lacks generalizability. Recent studies [weng2024digitaltwin, swaminathan24leia, deng2024articulate] have improved the accuracy of motion parameter prediction and demonstrated some capability in multi-part object reconstruction, but they struggle to generalize to objects with more parts and more complex motion combinations. ArtGS [liu2025artgs], built upon 3DGS, achieves certain improvements on multi-part objects, but its results are highly unstable and show poor generalization in large-scale evaluations. In this work, we propose a well-designed unified scalable reconstruction pipeline based on articulated Gaussians.
@@ 3.1 Preliminaries
3DGS [kerbl20233dgs] represents a 3D scene by a set of Gaussian Primitives, each defined as:
@@
To reconstruct articulated objects with multiple 1-DoF motion parts, we follow the two-state observation setting in DigitalTwinArt [weng2024digitaltwin], requiring posed RGB-D sequences of the same scene in two states as input: , , where denotes the number of images, the RGB image, the depth map, the camera extrinsics, and the camera intrinsics..
@@ 3.3 Soft-to-hard Training Paradigm
Sometimes, part segmentation alone cannot fully constrain motion parameter learning, making extreme movements and occlusions difficult to model. To refine motion learning, we apply a feature-matching-based trajectory regularization term during hard training. For image from view at state , we select the top closest view from state and apply an image feature matching model [sun2021loftr] to establish 2D pixel correspondences , where denotes the number of matches and are pixel coordinates. We lift them to 3D space as using depth and camera parameters. Then, we apply a 3D locality filter on 3D maches to mitigate false correspondences, yielding a refined set of matching pairs . The transformation at the matching positions can be derived from Eq. 6. The trajectory regularization term is the average difference between the position obtained through transformation and that derived from the filtered correspondences (Fig. 3(b)):
@@ 3.4 Optimization
where , is the D-SSIM loss [kerbl20233dgs], and . We also use the segmentation loss:
@@ 3.5 Initialization
Part segmentation. To initialize and regularize Gaussian part weights , we leverage SAM2 [ravi2024sam2], a foundation segmentation model pretrained on large-scale data. Since its zero-shot results often show inconsistent granularity on articulated objects, we fine-tune it on multi-view images and masks rendered from Partnet-Mobility (PM) [Xiang_2020_SAPIEN], yielding a specialized model, Art-SAM. We then apply cross-view propagation to ensure multi-view consistency (see Supplementary Material Secs. 9 and 10). This method is robust and generalizable, as Art-SAM requires only light post-training to adapt to novel object categories or configurations.
@@ 4.1 Benchmark Data Generation
Articulated objects are mainly constructed from the PM dataset [Xiang_2020_SAPIEN], from which we select 87 objects based on the diversity of categories, part number, and appearances. We use Blender with a procedural rendering pipeline [blender, Denninger2023blenderproc2] to render multi-view images of the 3D models in the base assets. To increase the diversity of object states while providing adequate observation of the interior parts of the objects, we set the 1-DoF part-level motion parameter to two random states: the starting state lies between and the ending state is within (here denotes the ”fully-closed” state and denotes the ”fully-open” state). For the object at each motion state, we place the camera in a spherical region around the object and randomly sample 100 views for training and 20 views for testing, all at a resolution of .
Due to the limited availability of high-quality real-world articulated object data and the significant ground-truth errors caused by the inability to annotate the internal structures of real objects, we selected only three well-annotated real objects from the Multiscan [mao2022multiscan] dataset. See Supplementary Material Sec. 18 for more results.
| 2 Parts (34) | 3 Parts (21) | 4-5 Parts (24) | 6-20 Parts (11) | All (90) | ||
| Axis Ang | ArtGS [liu2025artgs] | =20.74pt,depth=17.28pt,width=0.0pt | ||||
| Ours | =20.74pt,depth=17.28pt,width=0.0pt | |||||
| Axis Pos | ArtGS [liu2025artgs] | =20.74pt,depth=17.28pt,width=0.0pt | ||||
| Ours | =20.74pt,depth=17.28pt,width=0.0pt | |||||
| Part Motion | ArtGS [liu2025artgs] | =20.74pt,depth=17.28pt,width=0.0pt | ||||
| Ours | =20.74pt,depth=17.28pt,width=0.0pt | |||||
| CD-s | ArtGS [liu2025artgs] | =20.74pt,depth=17.28pt,width=0.0pt | ||||
| Ours | =20.74pt,depth=17.28pt,width=0.0pt | |||||
| CD-m | ArtGS [liu2025artgs] | =20.74pt,depth=17.28pt,width=0.0pt | ||||
| Ours | =20.74pt,depth=17.28pt,width=0.0pt |
@@ 5.1 Implementations
We first evaluate several methods on several traditional datasets, and finally select our method and ArtGS [liu2025artgs] for scale-up evaluation on our MPArt-90 benchmark.
@@ 5.3 Ablation Studies
Part assignment. We evaluate our model under 3 different scenarios: (1) without part segmentation masks, (2) without part initialization, and (3) replacing the part assignment module with an MLP, as used in GART [lei2024gart]. As shown in Table 2, without segmentation masks, Gaussians struggle to capture the complex motion of multiple parts. Additionally, relying solely on part segmentation supervision—without explicit part initialization—can negatively impact the geometric reconstruction of movable parts. When using MLPs for part assignment, the model fails entirely to learn part segmentation and motion for multi-part objects.
@@ 6.2 HSI
Furthermore, our Articulated Gaussians can be leveraged to generate 4D assets, enabling the modeling of HSI in dynamic environments. As depicted in Fig. 7, given Gaussian-based representations of humans, objects, and scenes as input, inspired by ZeroHSI [li2024zerohsi], we could generate high-fidelity 4D scenes. We first render a static frame, which is then used as input to a text-guided video diffusion model to synthesize a corresponding video. This generated video serves as a supervisory signal to optimize the kinematic parameters of humans and objects, effectively lifting the original 3D assets into a coherent 4D representation.
Specifically, we employ SMPL [loper2023smpl] for human kinematic modeling and use our reconstructed digital twins, with articulation parameters, to represent objects with kinematics. Through a distillation process, the motion dynamics captured in the video are transferred to the 4D scene.
@@ 9 Art-SAM Training
To provide supervision for accurate part-level segmentation, we adopt an image segmentation model to generate segmentation masks for each view and then use multi-view reprojection consistency to propagate monocular masks across views. To this end, an image segmentation model is required. Visual foundation models [kirillov2023segment, ravi2024sam2] have emerged as powerful tools for class-agnostic zero-shot or prompt-based image segmentation. However, when applied to articulated objects, the problems of over-segmentation and part ambiguity are common. Applying the models in a zero-shot manner often results in masks of erroneous granularity, where some masks cover overly fine-grained areas, while others fail to cover the full range of a complete interactable part (see the ’Zero-shot’ results of Fig. 10). Therefore, we fine-tuned SAM-v2.1 model [ravi2024sam2] to generate segmentation masks for articulated objects at the desired granularity. The fine-tuned model can generate masks of the desired part-level granularity on synthetic objects (Fig. 10) and real objects (Fig. 11). This section covers the details of data preparation and the fine-tuning process.
Base Assets. We use the PM dataset [Xiang_2020_SAPIEN] to create a dataset for fine-tuning SAM-v2.1. The dataset contains URDF-format models of articulated objects with motion parameters. We select 207 objects from 20 categories to generate pairs of images and corresponding segmentation masks for fine-tuning. For real-world assets, we select 50 objects from Multiscan [mao2022multiscan].
Image Rendering. We use Blender with a procedural rendering pipeline [blender, Denninger2023blenderproc2] to render multi-view images of the 3D models in the base assets. To increase the diversity of object states while providing adequate observation of the interior parts of the objects, we set the 1-DoF part-level motion parameter to random states between (here denotes the ”fully closed” state and denotes the ”fully open” state). For each object, we place the camera on a spherical region around the object and randomly sample 100 to 300 views, generating 130,000 training images at three resolution levels: , , and . Examples of training data are shown in Fig. 8.
Segmentation Map Rendering. To provide supervision at the desired granularity, we also adopt the procedural rendering pipeline with Blender [Denninger2023blenderproc2], which allows the rendering of object masks. A training sample consists of an RGB image paired with its corresponding segmentation masks.
Training Setup. Starting from the pretrained SAM2.1-Hiera-B+ model, we fine-tune all model parameters for 40 epochs on the dataset we created from PM. The loss function is a weighted linear combination of MSE IoU loss, Dice loss [milletari2016vnet], and focal loss [Lin_2017_Focal_Loss]. In each epoch, the learning rate starts at and decays to using a cosine annealing scheduler.